Suppose you are offered a project with the following payments:

Year Cash Flows 0 $ 8,900

1 − 4,600

2 − 3,300

3 − 2,400

4 − 1,700 a.

What is the IRR of this offer? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

IRR 15.96 15.96 Correct %

d-1. What is the NPV of the offer if the appropriate discount rate is 11 percent? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

NPV ???

d-2. What is the NPV of the offer if the appropriate discount rate is 23 percent? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

NPV ??
Answer #1

B8 fx =IRR(B3:37) A 1 Project A 2 year C ash flows pv@11% Present value 0 $ 8,900.00 8,900.00 1.0000 $ 1 $ (4,600.00)| 0.9009

B8 fc IRR(B3:37) 1 Project A 2 year 30 pv@11% 52 Cash flows 8900 -4600 - 3300 -2400 -1700 =IRR(B3:37) =C3/1.11 =C4/1.11 =C5/1